# Evolution Dynamics Toward the Limit Cycle of a Quantum Self-Sustained Oscillator

### v1

@@ Abstract
When subjected to a sufficiently weak, uni- or bi-directional coupling to one another, two or more oscillators may progressively match their rhythms to become synchronized—being under either frequency or phase entrainment [1, 2]. Synchronization is ubiquitous in nature—the circadian rhythm driven by external light [3, 4], flashing of bioluminescent insects [5, 6], a Belousov–Zhabotinsky reaction under light pulses [7], and a laser network [8], to name a few, and serves as a key mechanism in establishing order out of disorder in complex systems [9, 10, 11]. For synchronization to occur, an oscillator must have its own rhythm—the self-sustaining property is necessary for synchronization. A self-sustained oscillator is characterized by at least one source of energy dissipation (damping) and energy pump, such that there is a dynamical equilibrium between the opposing forces. This requires nonlinearities to be present in the system. In the phase space representation of the oscillator, the equilibrium corresponds to a limit cycle to which an arbitrary phase point is always attracted [1, 2]. The form of the limit cycle depends on the mathematical description of the system [12, 13]. Multiple limit cycles can exist for a more complex system [14]. Additional examples of self-sustaining oscillators include wall clocks, electronic systems [15], circadian rhythms [16, 17], chemical reactions [18, 19], and predator-and-prey systems [20].
@@
Figure 1: Wigner function evolution snapshots of the quantum RvdP oscillator for different initial states in the weakly nonlinear regime with $\kappa_{1}=0.1,\gamma_{1}=0.09,\gamma_{2}=0.001$. The same color value corresponds to different Wigner function values between the plots—this is done to emphasize the maximum of the Wigner functions. Negative Wigner function values (only at the central region of the top-left plot) are colored pale red. The snapshots showcase significant stages of the evolution toward the limit cycle, with the bottom row showing the Wigner functions reaching their steady states. The yellow circle is the classical limit cycle and the red point in the plots for the coherent state is the classical phase point at the given time. The initial states have $\left\langle{n}\right\rangle=1$. One simple harmonic period is $2\pi$ under the specified unit system.
@@
A coherent state evolves in a largely different manner, due to the presence of $m\neq 0$ off-diagonal elements and hence coherence decay, in addition to diagonal element redistribution. In this case, we observe a rotation of period $2\pi$ about the origin in the phase space, similar to the classical case, as the WF radially approaches the limit cycle radius. However, the classical phase point at any given time does not always coincide with the WF maximum as the evolution proceeds. This can be most easily seen in the third-row plot for the coherent state case in Fig. 1. We observe the WF smearing angularly in addition to smearing radially, which shows that the system gradually loses its phase. It is noteworthy that our observation here does not agree with the description of Ben Arosh et al. [21] that the Wigner function reaches the limit cycle first and then starts to lose its phase. We shall not describe the result for this case in terms of the OPD since the WF is also influenced by the $m\neq 0$ elements, making the relation between OPD and the form of the WF more complicated.
Figure 3: Time $T_{\mathrm{ss}}$ required to reach the steady-state limit cycle as a function of nonlinearity quantified by $\gamma_{2}/(\kappa_{1}-\gamma_{1})$, recorded as the trace distance to the steady-state density matrix falls below $10^{-3}$. Here $\kappa_{1}=0.1,\gamma_{1}=0.09$. Each plot shows the simulation result for initial states with different expected numbers $\left\langle{n}\right\rangle_{0}$ of excitation. In particular, the first and last points of the lines in the top left plot give the steady-state time for the evolution depicted in Figs. 1 and 2, respectively. The value for a given initial state with a given nonlinearity shows no particular trend with increasing $\left\langle{n}\right\rangle_{0}$—it may increase, decrease, or fluctuate.
@@
Figure 4: Calculated quantities to explain the long steady-state time in the coherent state case. The limit cycle is reached when all of the quantities presented are zero. The solid black line is the sum of the difference between the diagonal entries of the density matrix at time $t$ and the steady-state density matrix. The other lines are the sums of the absolute values of the $m$th off-diagonal entries of the density matrix at time $t$. A truncated Hilbert space of dimension $N=12$ is used, meaning that $|m|<12$. The rest of the sums are omitted as they start small and decay quickly. Calculated in the strongly nonlinear regime with $\kappa_{1}=0.1,\gamma_{1}=0.09,\gamma_{2}=0.1$, initialized as a coherent state with $\left\langle{n}\right\rangle_{0}=1$.

### v2

@@ Abstract
When subjected to a sufficiently weak, uni- or bi-directional coupling to one another, two or more oscillators may progressively match their rhythms to become synchronized—being under either frequency or phase entrainment [1, 2]. Synchronization is ubiquitous in nature—the circadian rhythm driven by external light [3, 4], flashing of bioluminescent insects [5, 6], a Belousov–Zhabotinsky reaction under light pulses [7], and a laser network [8], to name a few—and it serves as a key mechanism in establishing order out of disorder in complex systems [9, 10, 11]. For synchronization to occur, an oscillator must have its own rhythm—the self-sustaining property is necessary for synchronization. A self-sustained oscillator is characterized by at least one source of energy dissipation (damping) and energy pump, such that there is a dynamical equilibrium between the opposing forces. This requires nonlinearities to be present in the system. In the phase space representation of the oscillator, the equilibrium corresponds to a limit cycle to which an arbitrary phase point is always attracted [1, 2]. The form of the limit cycle depends on the mathematical description of the system [12, 13]. Multiple limit cycles can exist for a more complex system [14]. Additional examples of self-sustaining oscillators include wall clocks, electronic systems [15], circadian rhythms [16, 17], chemical reactions [18, 19], and predator-and-prey systems [20].
@@
Figure 1: Wigner function evolution snapshots of the quantum RvdP oscillator for different initial states in the weakly nonlinear regime with $\kappa_{1}=0.1,\gamma_{1}=0.09,\gamma_{2}=0.001$. The same color values correspond to different Wigner function values between the plots—this is done to emphasize the maximum of the Wigner functions. Negative Wigner function values (only at the central region of the top-left plot) are colored pale red. The snapshots showcase significant stages of the evolution toward the limit cycle, with the bottom row showing the Wigner functions reaching their steady states. The yellow circle is the classical limit cycle while the red point in the plots for the coherent state is the classical phase point at the given time. The initial states have $\left\langle{n}\right\rangle=1$. One simple harmonic period is $2\pi$ under the specified unit system.
@@
A coherent state evolves in a largely different manner, due to the presence of $m\neq 0$ off-diagonal elements and hence coherence decay, in addition to diagonal element redistribution. In this case, we observe a rotation of period $2\pi$ about the origin in the phase space at uniform speed, similar to the classical case, as the WF radially approaches the limit cycle radius. However, the classical phase point at any given time does not always coincide with the WF maximum as the evolution proceeds. This can be most easily seen in the third-row plot for the coherent state case in Fig. 1. We observe the WF smearing angularly in addition to smearing radially, which shows that the system gradually loses its phase. It is noteworthy that our observation here does not agree with the description of Ben Arosh et al. [21] that the Wigner function reaches the limit cycle first and then starts to lose its phase. We shall not describe the result for this case in terms of the OPD since the WF is also influenced by the $m\neq 0$ elements, making the relation between OPD and the form of the WF more complicated.
Figure 3: Time $T_{\mathrm{ss}}$ required to reach the steady-state limit cycle as a function of nonlinearity quantified by $\gamma_{2}/(\kappa_{1}-\gamma_{1})$, recorded as the trace distance to the steady-state density matrix falls below $10^{-3}$. Here $\kappa_{1}=0.1,\gamma_{1}=0.09$. Each plot shows the simulation result for initial states with different expected numbers $\left\langle{n}\right\rangle_{0}$ of excitation. In particular, the first and last points of the lines in the top left panel give the steady-state time for the evolution depicted in Figs. 1 and 2, respectively. The value for a given initial state with a given nonlinearity shows no particular trend with increasing $\left\langle{n}\right\rangle_{0}$—it may increase, decrease, or fluctuate.
@@
Figure 4: Distance between the $m$th off-diagonal from its steady state, as defined in Eq. (6). A truncated Hilbert space of dimension $N=12$ is used, meaning that $|m|<12$. The rest of the lines are omitted as they start small and decay quickly, cluttering the bottom left part of the plot. The calculation is performed in the strongly nonlinear regime with $\kappa_{1}=0.1,\gamma_{1}=0.09,\gamma_{2}=0.1$, initialized as a coherent state with $\left\langle{n}\right\rangle_{0}=1$.